OpenStudy (anonymous):

Using quadractic formula, please solve for me, x +11x-12=0

OpenStudy (anonymous):

it should be x2+11x-12=0

OpenStudy (anonymous):

u can solve this by factorising

OpenStudy (amistre64):

the quad formula is simply plug and play; why would you need someone to do it for you?

OpenStudy (anonymous):

gernic rectangels

OpenStudy (anonymous):

I wanna make sure i'm coming up with the right answer Amistre

OpenStudy (amistre64):

whats your answer then? and we can check it ;)

OpenStudy (anonymous):

-11 13squared

OpenStudy (anonymous):

oops

OpenStudy (amistre64):

it actually checks itself ; plug in the x values you find and see if they make it equal 0

OpenStudy (anonymous):

oh ok:/

OpenStudy (amistre64):

you should come up with x = -12 and 1, if i see it right

OpenStudy (anonymous):

yep amis

OpenStudy (amistre64):

-b sqrt(b^2 -4ac) -- +- ------------- 2a 2a

OpenStudy (amistre64):

-11 sqrt(121 +48) ---- +- ------------ 2 2

OpenStudy (amistre64):

-11 +- sqrt(169) -11 +- 13 --------------- = ---------- 2 2 -24/2 = -12 2/2 = 1

OpenStudy (anonymous):

Thanks Amis don't know where i went wrong

OpenStudy (amistre64):

you were close; you just forgot the "/2" and also to make sure that you use the "+-" and not just the a "+" for the sqrt

OpenStudy (anonymous):

Oh ok